OpenStudy (anonymous):

find the derivative of (3^2x)/x

OpenStudy (ranga):

(u/v)' = (vu' - u v') / v^2

OpenStudy (anonymous):

how do i set it up? find u and du?

rvc (rvc):

u=3^2x v=x

OpenStudy (accessdenied):

That is the general rule for taking the derivative of a quotient. \( \dfrac{d}{dx} \left( \dfrac{3^{2x}}{x} \right) = \dfrac{d}{dx} \left( \dfrac{u}{v} \right) \) Just compare the numerator and denominator. u = 3^2x and v = x. You see that better here?

OpenStudy (anonymous):

ah, so is it u sub or what do i do next

OpenStudy (accessdenied):

You'll need to compute the derivative of u and v. Then its a direct substitution into that rule formula.

OpenStudy (anonymous):

so du is 0 and dv=1???

OpenStudy (accessdenied):

Woah, careful. \( 3^{2x} \) is not a constant. Do you know the derivative of exponential functions?

rvc (rvc):

\[xdy/dx(3^2^{x}) - 3^2^{x}dy/dx(x)/x^2\]

rvc (rvc):

@pnp97

OpenStudy (accessdenied):

In a sense of exponential functions with a non-e base, you can make the adjustment with a logarithm property: \( N = e^{\ln N} \) \( 3^{2x} = \left(3^x\right)^2 = \left( e^{x \ \ln 3}\right)^2 \) \( \ln 3\) is a constant now. And the derivative of e^u is just itself. multiplied by the derivative of u.

OpenStudy (anonymous):

oh so once you adjust it to ln3 the derivative of that is itself so du=ln3

OpenStudy (anonymous):

@rvc is that it solved completely???

OpenStudy (accessdenied):

\( \dfrac{d}{dx} \left( 3^{2x} \right) = \dfrac{d}{dx} \left( e^{2 \ln 3 x} \right) = e^{2 \ln 3 x} * 2 \ln 3 \) Note chain rule comes in to play. This one is understandably tricky. So take your time and think about it. :)

OpenStudy (accessdenied):

\( \dfrac{d}{dx} \left( 3^{2x} \right) = \dfrac{d}{dx} \left( e^{2 \ln 3 \ \ x} \right) = e^{2 \ln 3 \ \ x} * 2 \ln 3 \) That x is not inside the argument of \( \ln 3\).

OpenStudy (accessdenied):

After that, you are correct in the d/dx ( x) = 1 So we would then just directly substitute those expressions we found: \( \dfrac{d}{dx} \left( \dfrac{u}{v} \right) =\dfrac{du/dx \ \ v - u \ \ dv/dx}{v^2} \)

OpenStudy (anonymous):

alright so du is e^2ln3x*2ln3????

OpenStudy (accessdenied):

Yes. (a small simplification is also to reverse that trick with the logarithm to make it look more like the u= 3^(2x) ==> \(du/dx = e^{2 \ln 3 \\ x} * 2 \ln 3 = 3^{2x} * 2 \ln 3\)

OpenStudy (anonymous):

ad do after that it would be d/dx (3^2x)/(x)) = (3^2x*2ln3)/dx (whats dx?) x-3^2x 1/dx)/1

OpenStudy (accessdenied):

\( \dfrac{d}{dx} \left( \dfrac{3^{2x}}{x} \right) = \dfrac{3^{2x} 2 \ln 3 \times x - 1 \times 3^{2x}}{x^2} \) I should emphasize that it is du/dx is the whole expression of the derivative of u with respect to x. du is just a differential and is equal to the derivative times dx.

OpenStudy (accessdenied):

Despite the variables being u and v (like an integration by parts), du/dx and dv/dx are the full expressions of derivatives. du/dx = u'(x). du = u'(x) dx.

OpenStudy (anonymous):

that makes sense! does the expression have to be simplied and combine the like terms i.e. 3^2x??/

OpenStudy (accessdenied):

for calculus, i know my teacher was lenient about it. as long as you had all the derivatives simplified out, you could leave it at that point. it might just depend on what your teacher considers though.

OpenStudy (anonymous):

i think it should be fine! thanks so much!

OpenStudy (accessdenied):

You're welcome! :)

rvc (rvc):

All the best @pnp97